### v1

# Adapting Large Multimodal Models to Distribution Shifts: The Role of In-Context Learning

Abstract
Recent studies indicate that large multimodal models (LMMs) are highly robust against natural distribution shifts, often surpassing previous baselines. Despite this, domain-specific adaptation is still necessary, particularly in specialized areas like healthcare. Due to the impracticality of fine-tuning LMMs given their vast parameter space, this work investigates in-context learning (ICL) as an effective alternative for enhancing LMMs’ adaptability. We find that the success of ICL heavily relies on the choice of demonstration, mirroring challenges seen in large language models but introducing unique complexities for LMMs facing distribution shifts. Our study addresses this by evaluating an unsupervised ICL method, TopKNearestPR, which selects in-context examples through a nearest example search based on feature similarity. We uncover that its effectiveness is limited by the deficiencies of pre-trained vision encoders under distribution shift scenarios, evidenced by their zero-shot capabilities barely outperforming random guesses. To address these challenges, we propose InvariantSelectPR, a novel method leveraging Class-conditioned Contrastive Invariance (CCI) for more robust demonstration selection. Specifically, CCI enhances pre-trained vision encoders by improving their discriminative capabilities across different classes and ensuring invariance to domain-specific variations. This enhancement allows the encoders to effectively identify and retrieve the most informative examples, which are then used to guide LMMs in adapting to new query samples under varying distributions. Our experiments show that InvariantSelectPR substantially improves the adaptability of LMMs, achieving significant performance gains on benchmark datasets, with a $34.2\%$$\uparrow$ accuracy increase in 7-shot on Camelyon17 and $16.9\%$$\uparrow$ increase in 7-shot on HAM10000 compared to the baseline zero-shot performance. Our code will be publicly available at: https://github.com/jameszhou-gl/icl-distribution-shift.
$\dagger$\$$\dagger$\$footnotetext: These authors contributed equally to this work.
Figure 1: Comparative illustration of (a) zero-shot transfer, which relies on LMMs’ pre-trained knowledge to respond to queries, potentially leading to a large distribution gap, and (b) in-context learning (ICL), which introduces an example from a closer distribution to bridge this gap. This work investigates different retrieval methods for selecting effective ICL examples. (c) The efficacy of one-shot ICL is showcased in guiding LMMs with shifted distributions.
1 Introduction
Machine learning models are essential in areas such as climate modeling, biomedicine, and autonomous driving, where they need to reliably manage deviations from their training data known as distribution shifts (Park et al., 2021; Zhou et al., 2022, 2023b). Traditional methods like domain adaptation and domain generalization have been somewhat effective but still fall short in addressing these shifts, as confirmed by several empirical studies (Gulrajani and Lopez-Paz, 2020; Wiles et al., 2022). However, the emergence of foundation models, characterized by their extensive and diverse pretraining, offers new possibilities for enhancing adaptability to these challenges (Bommasani et al., 2021; Radford et al., 2021; Shu et al., 2023). Specifically, large multimodal models (LMMs) (Yang et al., 2023b) such as GPT-4V (OpenAI, 2023), and Gemini (Team et al., 2023) have shown superior adaptability. Their zero-shot [^1] We adopt the zero-shot setting in the CLIP study (Radford et al., 2021), which enables zero-shot transfer via paired image-text features without extra linear probes. It differs from the classical zero-shot of generalizing to unseen categories (Chen et al., 2021, 2022, 2023). capabilities have been found to frequently outperform the performance of traditional fine-tuned models in natural datasets (Han et al., 2024).
Despite recent advances, domain-specific adaptation remains a significant challenge, especially in healthcare and scientific research (Han et al., 2024). While LMMs like Google DeepMind’s Med-Gemini offer fine-tuned versions for medical tasks (Saab et al., 2024), their block-box nature and massive parameter sets make traditional fine-tuning impractical for researchers without extensive computational resources. This highlights the urgent need for more feasible adaptation techniques. In-context learning (ICL), which allows models to adapt during inference without parameter adjustments, emerges as a promising alternative (Brown et al., 2020; Liu et al., 2021; Min et al., 2022; Dong et al., 2022). While the effectiveness of ICL is recognized within large language models (LLMs), its application for improving adaptability in LMMs under distribution shifts is less explored.
As depicted in Figure 1, we hypothesize that equipping LMMs with context examples that include task-specific information and details about the query sample can substantially enhance their performance. Our research starts with a thorough evaluation of ICL’s capacity to tailor LMMs (§2) to specific domains, particularly healthcare research, where there is a clear necessity for domain-specific fine-tuning (Han et al., 2024). We discover that the success of ICL heavily depends on the choice of demonstrations, supporting findings from previous LLMs studies (Liu et al., 2021; Min et al., 2022; Dong et al., 2022). Although the significant impact of demonstration selection on ICL performance is not entirely unexpected, the challenge of selecting demonstrations for LMMs under distribution shifts remains unexplored. To address this, we re-examine the unsupervised retrieval of in-context examples (§3.1), TopKNearestPR, traditionally used in LLMs. This method, discussed in (Liu et al., 2021; Zhang et al., 2024), uses feature similarity to pinpoint contextually relevant ICL examples.
However, the TopKNearestPR approach faces considerable challenges when applied to LMMs during distribution shifts. Notably, using pretrained vision encoders like CLIP-ViT [^2] https://huggingface.co/openai/clip-vit-base-patch16, zero-shot performance often remains at levels comparable to random guessing in specialized domains, as shown in a recent study (Han et al., 2024). This poor performance reveals a critical limitation in these encoders: they struggle to recognize and adapt to the subtle variations in new distributions, which compromises the reliability of visual feature similarities for selecting effective demonstrations. To tackle these challenges, we propose InvariantSelectPR, a novel method designed specifically for scenarios involving distribution shifts (§3.2). This approach employs Class-conditioned Contrastive Invariance (CCI) to choose demonstrations based on domain-invariant features, which are inherently robust to distributional changes (Zhou et al., 2024). This retriever method is distinctively crafted for distribution shifts, ensuring the resilience of selected in-context examples in varying conditions. Our empirical results demonstrate that InvariantSelectPR significantly improves the adaptability of LMMs, achieving notable accuracy improvements, i.e., a 34.2% accuracy improvement in 7-shot on Camelyon17 and a 16.9% accuracy increase in 7-shot on HAM10000 over the zero-shot baseline.
Our contributions and the key findings are summarized as follows: (1) To the best of our knowledge, this work takes the first step towards deeply understanding the role of in-context learning as an effective strategy for enhancing the adaptability of LMMs under distribution shifts (§2). (2) We introduce InvariantSelectPR, a novel in-context retrieval framework specifically developed to tackle distribution shifts (§3). (3) Through extensive experiments on four benchmark datasets (§4), our InvariantSelectPR method shows substantial enhancements over zero-shot generalization capabilities.
Figure 2: Motivation illustration: (a) Performance comparison between Zero-shot and RandomPR, illustrating the limitations of random in-context example selection across four datasets, where one-shot RandomPR often underperforms compared to zero-shot. (b) Analysis of 77 query samples from the target domain, hospital_3 in Camelyon17, using 50 distinct one-shot examples to examine performance variability. Mean values are marked in blue, and variance is represented by black lines, highlighting the significant impact of example selection on model accuracy. If appropriate in-context samples are chosen, there is a potential for gains up to 40.25%.
2 Motivation: ICL Demonstrations under Distribution Shifts
In this section, we evaluate ICL’s capability to enhance LMM adaptability. Starting with RandomPR, we randomly select in-context examples from source domain data without relevance to the target task. Our evaluation uses the Gemini model, noted for its zero-shot capabilities, across four medical datasets typically needing domain-specific fine-tuning (Han et al., 2024). We compare one-shot RandomPR with the zero-shot baseline for a preliminary investigation. According to Figure 2(a), while RandomPR presents a slight decrease of 0.3% on the Camelyon17 dataset, it leads to a substantial performance decline of 4.2%, 3.7%, and even 8.2% on the HAM10000, NIH_Chest, and COVID datasets respectively. Despite its conceptual simplicity, our empirical results suggest that random in-context example selection often fails to fulfill the essential requirement for effective model adaptation—providing informative and contextually appropriate demonstrations.
To unravel the variable efficacy of RandomPR, we conduct an experiment using the Camelyon17 dataset, focusing on 77 query samples from the target domain hospital_3. We test the influence of introducing 50 distinct examples from the source domains on the predictions for each query sample. The results in Figure 2(b), display both the mean and standard deviation, indicating significant variability in performance based on the in-context examples used. Notably, while zero-shot accuracy is 54.55% ($44/77$), our analysis reveals that up to 73 query samples could be accurately classified with the apt in-context samples, potentially boosting accuracy by 40.25%. Furthermore, the variability observed—such as a mean accuracy of 70% and a 21% variance for the 25th query—highlights the varying effects of different ICL examples. These findings highlight the inconsistencies in RandomPR’s performance and underscore the need for advanced methodologies in ICL example selection.
Figure 3: Overview of three retrieval methods: RandomPR, TopKNearestPR, and InvariantSelectPR. RandomPR selects examples without specific criteria, often overlooking informative ones. TopKNearestPR uses feature similarities for selection, yet struggles with domain-specific tasks where pre-trained encoder features lack sufficient detail. In contrast, InvariantSelectPR uses a class-conditioned contrastive invariance (CCI) framework to enhance vision encoders, effectively identifying the most representative samples by focusing on key invariant features.
3 Methodology
Upon identifying the limitations of RandomPR, we developed two advanced methods for more effective ICL example selection: TopKNearestPR and InvariantSelectPR, illustrated in Figure 3. These methods aim to enhance the adaptability of LMMs to distribution shifts through the strategic selection of demonstrative examples. We detail these selection methods below.
3.1 TopKNearestPR: Enhancing Context Relevance
TopKNearestPR adopts an unsupervised strategy to identify in-context examples by measuring the similarity between the feature vectors of a target query image $\bm{x}_{q}$ and those across $M$ source domains. The dataset $\mathcal{S}$ includes domains $\mathcal{S}^{i}=\{(\bm{x}_{j}^{i},\bm{y}_{j}^{i})\}_{j=1}^{n_{i}}$, where $\bm{x}$ represents feature vectors and $\bm{y}$ is class labels. The cosine similarity between the feature vectors from the query image $\bm{x}_{q}$ and any image $\bm{x}_{j}^{i}$ from the dataset, calculated using a pre-trained vision encoder like CLIP-ViT, is given by:
$\text{sim}(\bm{x}_{q},\bm{x}_{j}^{i})=\frac{\bm{z}(\bm{x}_{q})\cdot\bm{z}(\bm{x}_{j}^{i})}{\|\bm{z}(\bm{x}_{q})\|\|\bm{z}(\bm{x}_{j}^{i})\|}$ (1)
Here, $\bm{z}(\bm{x})$ refers to the feature vector extracted by the encoder. The top $K$ images that exhibit the highest similarity to the query are selected using:
$\text{top}_{K}\left(\{\text{sim}(\bm{x}_{q},\bm{x}_{j}^{i}):i=1,\ldots,M;j=1,\ldots,n_{i}\}\right)$ (2)
where $\text{top}_{K}$ denotes the operation of selecting the indices of the $K$ largest values from the set. The selected images serve as the in-context examples for the LMMs, aiming to enhance their understanding and performance on analogous tasks without further training.
3.2 InvariantSelectPR: Tailored for Distribution Shift Adaptation
TopKNearestPR focuses on relevance by utilizing feature similarities, but its effectiveness can be constrained by the granularity of features from conventional encoders. Pretrained vision encoders, such as CLIP-ViT, while robust in general scenarios, often struggle to differentiate effectively in domain-specific tasks. This limitation manifests as zero-shot performances that are only marginally better than random guesses, leading to the selection of suboptimal in-context examples when relying solely on pre-trained models. Thus, we propose InvariantSelectPR, a new method designed to enhance robustness across distribution shifts.
Facilitating Class-conditioned Contrastive Invariance.  InvariantSelectPR is centered around the Class-conditioned Contrastive Invariance (CCI) mechanism, which aims to improve the model’s ability to distinguish between classes while maintaining stability across domain-specific variations (Zhou et al., 2024). This is achieved by promoting similarity among instances of the same class from different domains and highlighting differences between classes. Using the class token embedding [CLS], $\bm{x}_{N}$, from the final vision transformer (ViT) layer, the CCI loss is defined as:
$\mathcal{L}_{\text{CCI}}=-\mathbb{E}\left[\log\frac{\exp(\bm{z}_{N}\cdot\bm{z}_{N^{\prime}}/\tau)}{\sum_{k\neq N}\exp(\bm{z}_{N}\cdot\bm{z}_{k}/\tau)}\right]$ (3)
Here, $\bm{z}_{N^{\prime}}$ is a positive sample of $\bm{z}_{N}$ from the same class but possibly a different domain, and $\bm{z}_{k}$ signifies a negative sample of $\bm{z}_{N}$ from a different class. $\tau$ denotes the temperature parameter in contrastive learning (Chen et al., 2020; Zhou et al., 2023a). This formulation ensures that the learned representations are both discriminative and invariant, crucial for adapting to new distributions.
This approach combines this CCI loss $\mathcal{L}_{\text{CCI}}$ with a classification loss $\mathcal{L}_{\text{cls}}$ to enhance the vision encoder’s ability to manage distribution shifts effectively. The classification loss uses cross-entropy to align the final class token embedding $\mathbf{x}_{N}$ with the ground-truth label $\bm{y}$, bolstering the model’s discriminative power:
$\mathcal{L}_{\text{cls}}=-\sum_{i=1}^{C}\bm{y}_{i}\log(\text{Head}(\bm{x}_{N})_{i})$ (4)
$\mathcal{L}_{\text{total}}=\mathcal{L}_{\text{cls}}+\lambda\mathcal{L}_{\text{CCI}}$ (5)
where $C$ represents the total number of classes in the dataset, and $\text{Head}(\cdot)$ is a neural classification head that maps the class token $\bm{x}_{N}$ to a predicted class probability distribution. $\lambda$ is a tuning hyper-parameter to control the weight of the CCI loss.
In-Context Selection Through Enhanced Invariance.  After fine-tuning the vision encoder with the combined losses in Eq. (5), we leverage refined features to assess the similarity between the target samples and in-context examples. By ensuring these similarities reflect both visual resemblance and domain invariance, the k-shot examples with the highest similarity scores are then selected.
4 Experiments
Datasets Overview. We use four benchmark datasets to explore distribution shifts, particularly emphasizing domain-specific fine-tuning (Han et al., 2024). Camelyon17 (Bandi et al., 2018) features 450,000 patches from breast cancer images across five hospitals. HAM10000 (Tschandl et al., 2018) offers dermatoscopic images critical for skin cancer detection. The NIH_Chest dataset (Wang et al., 2017) includes over 112,000 X-ray images annotated for thoracic diseases. The COVID dataset (Han et al., 2021) provides diverse pneumonia detection data, including COVID-19 cases, from various hospitals. We analyze a practical subset, random_1, with 450 samples [^3] Available at https://github.com/jameszhou-gl/gpt-4v-distribution-shift.
Implementation Details. We compare three retrieval methods—RandomPR, TopKNearestPR and InvariantSelectPR—against the baseline zero-shot capability. We employ vit_large_patch14_224_clip_laion2b configuration from the timm library, exploring variations in backbone configurations further in §4.3.2. The Gemini model is employed as the primary LMM due to its superior zero-shot performance across varied datasets (Han et al., 2024) and its stable log-linear improvement in performance with an increasing number of ICL examples, as observed in a concurrent study (Jiang et al., 2024). Our main results (§4.1) focus on one-shot performance, with additional insights on the impact of different numbers of shots in §4.3.3. We also include other leading LMMs, such as GPT-4V and Claude (Anthropic, 2023), in our extended analysis in §4.3.4.
[…]
4.1 Main Results
In Tables 1 and 2, our analysis of four benchmark datasets provides a detailed examination of how different ICL methods perform under distribution shifts. The zero-shot approach highlights the inherent ability of LMMs to adapt to new domains without retraining. However, the effectiveness of this adaptability varies significantly with different ICL strategies. The RandomPR strategy, which employs a stochastic method for selecting in-context examples, yields inconsistent results. For instance, on the Camelyon17 dataset, it leads to a slight decrease of 0.3% in accuracy, but it largely underperforms on the COVID, HAM10000, and NIH_Chest datasets, with accuracy decreases of 8.2%, 4.2%, and 3.7%, respectively. This highlights the unpredictable performance of RandomPR across different conditions. Conversely, TopKNearestPR uses a pre-trained vision encoder to identify feature similarities for example selection, leading to a 5.74% improvement on the Camelyon17, which demonstrates the benefits of a more targeted approach in example selection. Despite this success, the method sees declines of 3.9% and 2.9% on the HAM10000 and NIH_Chest datasets, respectively, indicating a lack of consistent performance across all test scenarios. The most effective strategy, InvariantSelectPR, consistently outperforms other methods, significantly exceeding the zero-shot baseline across all datasets, especially achieving remarkable gains of 8.3% on HAM10000 and 8.6% on Camelyon17. These results underscore the importance of advanced in-context example selection techniques in adapting LMMs to distribution shifts. Despite notable gains, the improvements with InvariantSelectPR on NIH_Chest and COVID are modest. In Figure 4, the incremental improvements by InvariantSelectPR align with those from fine-tuned encoders, which generally surpass the fine-tuning approach by 1% to 6%. This suggests that when fine-tuning itself is minimally effective, ICL strategies yield limited enhancements. Future research thus focuses on more sophisticated methods to enhance invariance, beyond fundamental domain-invariance in this work.
Table 1: Performance comparison of three retrieval methods against the zero-shot approach, illustrating accuracy improvements or decreases on Camelyon17 and COVID datasets. Mean and standard deviation values are calculated over three independent runs, with the best results highlighted in bold.
Method Camelyon17 COVID
Hosp0 Hosp1 Hosp2 Hosp3 Hosp4 Acc Sou Tar Acc
Zero-shot 52.00 51.93 56.44 54.55 56.67 54.17$\pm$0.5 62.19 44.19 52.75$\pm$1.3
RandomPR 50.50 53.03 54.59 53.28 58.55 53.87$\pm$1.1 38.66 49.86 44.52$\pm$0.9
TopKNearestPR 62.24 58.65 58.62 59.65 60.15 59.91$\pm$1.8 41.59 60.88 51.70$\pm$2.7
InvariantSelectPR 60.12 63.96 62.68 63.39 64.36 62.77$\pm$1.1 39.94 67.05 54.15$\pm$1.0
Table 2: Performance comparison of three retrieval methods against the zero-shot approach, illustrating accuracy improvements or decreases on HAM10000 and NIH_Chest datasets. Mean and standard deviation values are calculated over three independent runs, with the best results highlighted in bold.
Method HAM10000 NIH_Chest
RD VMod VMol VDis Acc PA AP Acc
Zero-shot 28.54 37.39 26.42 42.22 32.62$\pm$1.2 12.41 14.26 13.31$\pm$0.7
RandomPR 21.48 32.81 27.88 12.64 25.71±$\pm 1.9$ 7.91 10.12 8.98$\pm$2.1
TopKNearestPR 23.03 33.39 31.73 28.03 28.72$\pm$1.0 10.34 10.43 10.39$\pm$1.0
InvariantSelectPR 38.20 49.43 38.96 25.19 40.91$\pm$1.0 13.22 13.63 13.42$\pm$0.7
4.2 Ablation Study
To assess the impact of enhanced invariance on model adaptability, we conduct an ablation study focusing on the Gemini model’s one-shot performance. This study compares three configurations: a baseline using TopKNearestPR, the baseline only with $\mathcal{L}_{\text{cls}}$, and the full InvariantSelectPR that incorporates both $\mathcal{L}_{\text{cls}}$ and $\mathcal{L}_{\text{CCI}}$. Table 3 displays incremental performance gains across datasets with the successive additions of $\mathcal{L}_{\text{cls}}$ and $\mathcal{L}_{\text{CCI}}$. The addition of $\mathcal{L}_{\text{cls}}$ alone leads to a modest increase in performance by $1.77\%$. However, when incorporated with $\mathcal{L}_{\text{CCI}}$, there is a more substantial performance boost of $4.01\%$, confirming the effectiveness of CCI loss in improving the models’ adaptability.
Table 3: Ablation study on loss terms. The baseline is TopKNearestPR.
Configurations Camelyon17 COVID HAM10000 NIH_Chest Average
baseline 61.96 54.22 29.84 10.49 39.13
baseline+$\mathcal{L}_{cls}$ (w/o CCI) 61.59 52.00 38.93 11.11 40.90
baseline+$\mathcal{L}_{cls}$+$\mathcal{L}_{CCI}$ (full) 63.90 54.44 41.56 12.67 43.14
4.3 In-depth Analysis
4.3.1 ICL vs. Traditional Supervised Finetuning (SFT)
Recent advances in LMMs demonstrate their impressive zero-shot generalization, often outperforming fine-tuned models in natural distribution shifts (Han et al., 2024). This raises questions about whether LMMs with in-context learning, can exceed fine-tuned model performance in scientific datasets, traditionally reliant on domain-specific fine-tuning. We assess our ICL strategy against traditional SFT to explore this. For SFT, we focus on maintaining domain invariance and targeting the class prediction objective, similar to Eq. (5). We fine-tune a CLIP-ViT on source domain data and then apply it to predict outcomes on target examples. This comparison directly measures the effectiveness of ICL versus conventional SFT. For InvariantSelectPR, we choose ICL examples ranging from one to seven and report the best accuracy. Figure 4 displays the comparative performance across four datasets. SFT demonstrates a substantial improvement over zero-shot capabilities with accuracy improvements of 32.4%, and 12.3% on Camelyon17 and HAM10000, but underperforms 1.5% and 5.3% on NIH_Chest and COVID. In contrast, our proposed InvariantSelectPR with few-shot examples, consistently exceeds SFT, with gains of 1.4%, 4.4%, 1.6%, and 6.4% in the same datasets.
Figure 4: Comparative accuracies between Zero-shot, Supervised Finetuning, and InvariantSelectPR methods across various datasets, illustrating the superior performance of InvariantSelectPR over both zero-shot and supervised fine-tuning.
4.3.2 Backbone Evaluation
We evaluate the impact of different vision encoder backbones on the effectiveness of our InvariantSelectPR method compared to TopKNearestPR. This includes ViT models pretrained on ImageNet-21K and trained with the self-supervised DINO method on ImageNet-1K. Table 4 shows that InvariantSelectPR consistently outperforms TopKNearestPR across datasets, with an average accuracy of 42.9% versus 39.7%. This underscores the limitations of relying solely on pretrained visual similarity for selecting meaningful in-context examples. InvariantSelectPR also demonstrates more consistent performance, with less deviation from mean accuracy (under 1%) compared to TopKNearestPR (nearly 2%). An important observation is the enhanced performance of backbones utilizing self-supervised or contrastive learning methods, supporting the effectiveness of self-supervised learning in capturing generalizable features that contribute to more robust ICL performance, as suggested in studies (Chen et al., 2020; Radford et al., 2021; Wang et al., 2023a).
Table 4: Performance comparison of ICL methods with different vision encoder backbones.
Methods Backbones Camelyon17 COVID HAM10000 NIH_Chest Average
TopKNearestPR vit-l/14-clip 61.96 54.22 29.84 10.49 39.13
vit-l/16-in21k 60.45 49.78 30.80 10.24 37.82
vit-b/16-dino 63.72 59.68 32.74 10.89 41.76
InvariantSelectPR vit-l/14-clip 63.90 54.44 41.56 12.67 43.14
vit-l/16-in21k 61.76 52.78 38.15 12.89 41.40
vit-b/16-dino 64.48 53.72 44.55 11.36 43.53
4.3.3 ICL Examples with Various Shots
To assess the impact of the number of ICL examples, we perform an empirical study using the Camelyon17 and HAM10000 datasets, varying the number of shots from 1 to 7 for each dataset in Figure 5. This analysis reveals that increasing the number of shots leads to a decrease in the performance of the RandomPR method, implying that additional examples might introduce unhelpful information. In contrast, the TopKNearestPR method typically improves with more shots but shows a decline in performance when moving from 3-shot to 5-shot on the HAM10000 dataset, suggesting potential issues with example selection or redundancy. On the other hand, our InvariantSelectPR method consistently improves performance as the number of shots increases, demonstrating its effectiveness in utilizing information from source domains. Notably, this method achieves a performance boost of approximately 24.6% when the shot count increases from 1 to 7 on Camelyon17.
Figure 5: Performance comparison with varying numbers of ICL examples (shots) on Camelyon17 and HAM10000 datasets.
4.3.4 Evaluation Across Different LMMs
The open-source LMMs like IDEFICS (Laurençon et al., 2024) and OpenFlamingo (Awadalla et al., 2023) primarily focus on text and ignore the input signal of images (Bertini Baldassini et al., 2024). Furthermore, these LMMs lack instruction-following ability to choose the response from the answer list. Thus, we use three proprietary LMMs in this comparative analysis: Gemini Pro, GPT-4V, and Claude 3 Opus (Anthropic, 2024). Due to the high computational demands and associated costs of GPT-4V and Claude 3 Opus, we limit our testing to a single dataset, HAM10000, and perform a one-shot evaluation. Figure 6 demonstrates that InvariantSelectPR consistently outperforms other methods across all three LMMs. This method not only exceeds baseline zero-shot performance but also significantly enhances adaptability. Both GPT-4V and Claude 3 Opus exhibit substantial improvements using all ICL methods over their zero-shot capabilities, suggesting that ICL can effectively boost the adaptability of LMMs. This analysis highlights the capacity of InvariantSelectPR to leverage domain-invariant features to enhance LMMs performance under variable conditions.
Figure 6: Performance comparison of Zero-shot, RandomPR, TopKNearestPR, and InvariantSelectPR on the HAM10000 dataset across three LMMs, demonstrating the impact of one-shot demonstrations.
4.3.5 Distance Metric Evaluation
We examine the effects of employing various distance metrics, including Cosine, Euclidean, and Manhattan, within TopKNearestPR and InvariantSelectPR methods, as illustrated in Table 5. Our findings indicate that the InvariantSelectPR method consistently achieves higher performance than TopKNearestPR across all metrics tested on both the Camelyon17 and HAM10000 datasets.
Table 5: Performance comparison using different distance metrics.
Method Camelyon17 HAM10000
Cosine Euclidean Manhattan Avg Cosine Euclidean Manhattan Avg
TopKNearestPR 61.96 60.77 59.18 60.64 29.84 29.46 30.80 30.03
InvariantSelectPR 63.90 61.09 62.70 62.56 41.56 37.22 38.93 39.24
4.3.6 Computational Efficiency
Table 6 illustrates the trade-off between computational cost and accuracy improvement. While InvariantSelectPR incurs a slightly higher inference time and GPU usage than the zero-shot baseline but offers an 8.60% accuracy improvement. The increased cost is due to the model loading and similarity calculation. InvariantSelectPR’ lower inference time compared to TopKNearestPR is because it loads the vision encoder once per environment instead of for each target sample. Future work will focus on optimizing these steps to reduce inference time while maintaining accuracy gains.
Table 6: Performance comparison of different one-shot ICL methods on the Camelyon17 dataset, in terms of inference time, GPU usage, and accuracy improvement over the zero-shot baseline.
Method Inference Time (s/query) GPU Usage (GB) Accuracy Improvement
Zero-shot 5.23 - -
RandomPR 5.35 - -0.30%
TopKNearestPR 15.52 2.41 +5.74%
InvariantSelectPR 11.79 3.55 +8.60%
5 Conclusion and Discussion
Concusion. We investigated the efficacy of in-context learning (ICL) to improve the adaptability of LMMs to distribution shifts through our novel ICL approach, InvariantSelectPR. This method not only outperforms standard zero-shot capabilities but also exceeds other methods like RandomPR and TopKNearestPR in handling domain-specific shifts. Evaluations across four datasets confirmed that InvariantSelectPR enhances LMM adaptability by optimally selecting demonstrative examples. Our study offers insights for future work on distribution shifts in foundation models.
Limitations.  Our study is constrained by several factors. Firstly, we confined our analysis to a small selection of benchmark datasets and relied exclusively on commercial and proprietary models, such as Gemini Pro, GPT-4V, and Claude 3 Opus. The limited availability of comprehensive documentation for these models constrains our understanding of their pre-training data, architecture, and inherent biases. This is critical as some broadly used open-source LMMs can not effectively understand multiple images like Flamingo (Alayrac et al., 2022) and simultaneously follow instructions like LLaVA (Liu et al., 2023), necessitating the use of commercial models. Additionally, the substantial financial and computational resources required to access these proprietary models may restrict further validation and analysis. Secondly, our empirical tests involved just 450 samples, which, despite prior research suggesting stability ranging from 180 to 1800 cases (Han et al., 2024), might not reveal scalability issues or subtle biases in larger datasets. Thirdly, the prevalence of numerous domains in healthcare (Yang et al., 2023a) and scientific research (Ji et al., 2022) presents potential challenges in scaling our method across multiple domains.
Boarder Impacts.  This work introduces a novel ICL method to enhance the adaptability of LMMs to distribution shifts, particularly in scientific and healthcare domains. While this method promises improved model accuracy and reliability, its misuse could amplify biases or yield unreliable outputs from LMMs, leading to negative consequences in these critical fields.
Appendix
More contents are put in the appendix, including:
  • A.
    Related Work.
  • B.
    Experimental Details.
  • C.
    t-SNE Visualizations of Visual Features.
Appendix A Related Work
Distribution Shifts.  The literature on distribution shifts categorizes mitigation approaches into two primary strategies: domain adaptation and domain generalization. Domain adaptation techniques, well-established for scenarios where the target domain is known during training, recalibrate models according to the target data’s statistical properties [Ben-David et al., 2006]. These techniques encompass deep transfer learning, which aligns feature distributions between source and target domains [Sun and Saenko, 2016], unsupervised methods that minimize domain discrepancies [Ganin and Lempitsky, 2015], and the use of benchmarks such as Office-Home [Venkateswara et al., 2017] and DomainNet [Peng et al., 2019]. These benchmarks have propelled advances by introducing more complex and diverse scenarios [Han et al., 2022a, b]. In contrast, domain generalization addresses the more daunting challenge of excelling in completely unseen domains. Strategies here include aligning features across multiple source domains [Li et al., 2018b], separating domain-specific from domain-general features [Piratla et al., 2020, Chattopadhyay et al., 2020], employing meta-learning for optimization across various domains [Li et al., 2018a, Balaji et al., 2018], and using data augmentation to mimic domain variability [Volpi et al., 2018, Carlucci et al., 2019]. Recent studies have observed that LLMs have demonstrated exceptional adaptability when dealing with natural distribution shifts but cannot handle the distribution shifts in specialized areas such as healthcare and scientific research [Radford et al., 2021, Han et al., 2023]. This observation motivates this paper’s exploration of ICL and the development of new ICL strategies under distribution shifts.
In-Context Learning.  In-context learning (ICL), particularly defined in GPT-3 [Brown et al., 2020], originated in LLMs for natural language processing (NLP) tasks. ICL is a proven effective paradigm that leverages context augmented with a few examples to enable LLMs to make predictions [Dong et al., 2022, Lu et al., 2021, Wei et al., 2022b, Min et al., 2022, Dong et al., 2022, Wei et al., 2022a, Wolf et al., 2023, Wies et al., 2024, Xie et al., 2021]. The choice of these in-context examples critically impacts performance, as evidenced by studies demonstrating that selecting nearest neighbors based on sentence encoders can significantly enhance the few-shot capabilities of models like GPT-3 [Liu et al., 2021, Wu et al., 2024, Mao and Yu, 2024]. While ICL is established in NLP, it is emerging in visual and multimodal LLMs. The study Flamingo [Alayrac et al., 2022] marks the earliest exploration of visual ICL, with subsequent studies validating the importance of example selection in image painting models [Bar et al., 2022, Wang et al., 2023b, Zhang et al., 2024], visual understanding [Balazevic et al., 2024, Fang et al., 2024], and diffusion models [Wang et al., 2024]. Unlike prior work, this paper uniquely focuses on deeply understanding the role of ICL under distribution shifts, taking a first step in this direction.
Appendix B Experimental Details
B.1 Datasets
This section provides detailed information about the datasets used in the experiments, including their statistics, preprocessing steps, and domain splits.
Dataset Statistics.  We present the data statistics in Table 7. Besides, we use the datasets available in https://huggingface.co/datasets/jameszhou-gl/gpt-4v-distribution-shift and evaluate random_1 subset with 450 cases.
Table 7: Detailed statistics of the datasets used in the experiments.
Dataset Category Prediction Task Domain Type # Domains # Classes # Samples Example Image
Camelyon17 Medical Tumor detection Hospital 5 2 450 [Uncaptioned image]
HAM10000 Medical Skin disease classification Hospital 4 7 450 [Uncaptioned image]
NIH_Chest Medical Lung disease diagnosis Hospital 2 15 450 [Uncaptioned image]
COVID Medical Pneumonia type classification Hospital 2 3 450 [Uncaptioned image]
Preprocessing Steps.  We do not adopt additional preprocessing steps for the images, adhering instead to the guidelines provided by each large multimodal model. For instance, the Gemini model accepts PIL.Image.open(img_path) as input without requiring specific image size, normalization, or augmentation [^4] https://ai.google.dev/gemini-api/docs/get-started/python.
B.2 Prompts
We utilize the following basic prompt template in all ICL experiments, refering to [Han et al., 2024].
image_descriptions = [f"Image {i+1} is {image_class}" for i,
                      (desc, image_class) in enumerate(source_images)]
images_description = ". ".join(image_descriptions)
prompt = f"""Given the images, answer the following question, using the
         specified format.
         {images_description}.
         Question: What is the class of the next image?
         Choices: {’, ’.join(class_names)}.
         Please respond with the following format for each image:
         ---BEGIN FORMAT TEMPLATE---
         Answer Choice: [Your Answer Choice Here]
         Confidence Score: [Your Numerical Prediction Confidence Score Here
         From 0 To 1]
         Reasoning: [Your Reasoning Behind This Answer Here]
         ---END FORMAT TEMPLATE---
         Do not deviate from the above format.
         Repeat the format template for the answer.
B.3 Implementation Details
Vision encoder architectures.  We employ the vit_large_patch14_224_clip_laion2b configuration from the timm library [^5] https://github.com/huggingface/pytorch-image-models/blob/main/timm/models/vision_transformer.py, which utilizes the Vision Transformer (ViT) architecture [Dosovitskiy et al., 2020]. This model divides 224x224 pixel images into 14x14 patches and processes them using a transformer encoder, leveraging extensive pretraining on the LAION-2B dataset with the CLIP approach to enhance visual and textual understanding.
Data and Code.  Our code can be found at the supplementary material and data is publicly available in https://huggingface.co/datasets/jameszhou-gl/gpt-4v-distribution-shift.
Appendix C t-SNE Visualizations of Visual Features
In this section, we presnet t-SNE [Van der Maaten and Hinton, 2008] visualizations of the visual features to illustrate how class-conditioned contrastive invariance (CCI) contributes to domain invariance and discriminative capabilities. The visualizations are based on the original vision encoder and our fine-tuned vision encoder on three datasets, as shown in Figure 7. The visual features are extracted using both the pretrained and fine-tuned ViT models. The t-SNE plots are created to highlight the clustering behavior of the features from the target domain. By comparing the plots, we can visually assess the impact of the fine-tuning with CCI on the separation of different classes and the compactness of feature clusters.
Figure 7: t-SNE visualizations of visual features from the target domain for three datasets. The lower row shows features extracted using the pretrained ViT model, while the upper row shows features extracted using the our fine-tuned ViT model.
These t-SNE visualizations, particularly for the Camelyon17 dataset, clearly demonstrate that fine-tuning with class-conditioned contrastive invariance (CCI) significantly enhances the model’s ability to generalize across unseen domains. The fine-tuning process improves discriminative power by better aligning feature representations with class labels. This visualization underscores the critical importance of incorporating CCI to refine the vision encoder. By enhancing the alignment of feature representations with their corresponding classes, CCI contributes to a more robust and domain-invariant model. Furthermore, this refined vision encoder facilitates the selection of in-context learning (ICL) examples, enabling large multimodal models to adapt more effectively.